### v1

@@ 1 Introduction
The analysis of the X-ray radiation emitted from the strong gravity region around black holes is a powerful tool to study the accretion process of the innermost region of the accretion disk, measure black hole masses and spins, and test Einstein’s theory of General Relativity in the strong field regime. The prototype of astrophysical system for X-ray observations is shown in the left panel in Fig. 1 and is normally referred to as the disk-corona model [6, 7]. The central black hole can be either a stellar-mass black hole in an X-ray binary system or a supermassive black hole in an active galactic nucleus. The crucial ingredient is that the accretion disk is geometrically thin and optically thick, which requires a high value of the angular momentum of the accreting material and an accretion rate between $\sim 3$% and $\sim 30$% of the Eddington limit of the source. Such an accretion disk is cold because it can efficiently emits radiation. The thermal spectrum of the accretion disk turns out to be peaked in the soft X-ray band for stellar-mass black holes in X-ray binary systems and in the UV band for supermassive black holes in active galactic nuclei [8, 9]. The “corona” is some hot plasma ($\sim 100$ keV) near the black hole and the inner part of the accretion disk, but its exact morphology is not yet completely understood [10, 11, 12, 13, 14, 15, 16]. The corona may be the hot accreting material in the plunging region between the inner edge of the accretion disk and the black hole, the hot atmosphere above the accretion disk, the base of the jet, etc. (see the right panel in Fig. 1).
@@
    Are astrophysical black holes the Kerr black holes predicted by General Relativity? Do astrophysical black holes have an event horizon or are they instead compact horizonless objects as predicted by a number of scenarios beyond General Relativity [27]?
@@
    What is the mechanism responsible for QPOs? Are they generated in the accretion disk or in the corona? Can we use QPO observations for measuring black hole masses and spins and for testing Einstein’s theory of General Relativity?
@@
Figure 2: The top two panels show a simulated SFA light curve and spectrum of GRS 1915$+$105, demonstrating eXTP’s capabilities in studying strong-gravity accretion physics. The rapid, periodic variability reveals instabilities of the inner accretion disk of this black hole; the relativistic Fe K emission line provides insights into strong-gravity effects and accretion disk properties; and the absorption features from disk winds enable kinematic studies and an investigation of potential driving mechanisms. These topics will be further explored throughout this paper. The bottom two panels illustrate the expected number of black hole X-ray novae eXTP will observe over the first three-year period, based on historical events. These include short outbursts lasting a few months and longer outbursts extending over several years. See text for further details.
The enhanced X-ray Timing and Polarimetry (eXTP) mission has been proposed by a consortium led by the Institute of High-Energy Physics of the Chinese Academy of Sciences to address these questions. Its launch is currently scheduled in 2030. The primary goals of eXTP are: $i)$ the study of matter at very high density [28], $ii)$ the study of strong gravity with black holes (this paper), $iii)$ the study of matter in strong magnetic fields [29], $iv)$ the study of violent high-energy astrophysical phenomena with X-ray in the era of multi-messanger astrophysics [30] and $v$ the observatory science [31].
@@
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of $4\times 10^{-7}\,{\rm ergs\,cm^{-2}\,s^{-1}}$ (1 s exposure) across the 10–600 keV energy range, with an angular resolution of $20^{\prime}$, a position accuracy of 5’, and an energy resolution better than $30\%$ at 60 keV.
@@
Beyond all-sky monitoring, eXTP will conduct detailed studies of X-ray novae in the future, with a particular focus on black hole X-ray binaries in this paper. The top panels of Fig. 2 showcase a simulated SFA light curve and spectrum of the black hole X-ray binary GRS 1915+105 [39], based on archival observations from Insight-HXMT. eXTP will reveal rapid and periodic variability on short timescales associated with disk instability [see Section 2.2 or reference 40]. A 20 ks eXTP observation will also demonstrate the spectral capabilities of SFA, detecting relativistic Fe K disk emission as part of disk reflection (discussed in Section 2.2.1), which will provide insight into the properties of the strong gravity region and the accretion disk. Additionally, eXTP will resolve the absorption features of the disk winds, allowing measurements of their velocities and helping to determine their kinematic properties and potential launching mechanisms (see Section 3.1.2).
The paper is organized as follows. In Section 2, we show how eXTP can improve current measurements of black hole masses and spins as well as put stronger constraints on possible deviations from the predictions of General Relativity. In Section 3, we show how eXTP can help us to understand better the accretion process in the strong gravity region of black holes. Summary and conclusions are reported in Section 4.
2 Measuring the Basic Parameters of Black Holes and Testing General Relativity
@@
For a Kerr black hole, the mass $M$ is a free parameter, and therefore we may have black holes with arbitrarily low as well as arbitrarily high masses. However, the black holes in our Universe must form through some astrophysical process and therefore it is natural to expect black holes with masses only in specific ranges. This is exactly what we find. Astrophysical observations have discovered at least two classes of black holes: stellar-mass black holes ($M\approx 3-100$ $M_{\odot}$[45] and supermassive black holes ($M\sim 10^{5}-10^{10}$ $M_{\odot}$[46]. There is likely a third class of objects, the so-called intermediate mass black hole candidates [47], which are black hole candidates with a mass filling the gap between the stellar-mass and the supermassive ones.
@@
Unlike the mass, the spin angular momentum $J$ of a Kerr black hole is subject to certain constraints. If we define the dimensionless spin parameter $a_{*}=J/M^{2}$ (in units in which $c=G_{\rm N}=1$), the condition for the existence of a Kerr black hole is $|a_{*}|\leq 1$ (Kerr bound). For $|a_{*}|>1$, there is no event horizon and the Kerr solution describes the spacetime of a naked singularity, which is thought to be impossible to create in the Universe [48]. Spin measurements are crucial to understand how black holes form, evolve, and interact with the host environment [49]. In the case of stellar-mass black holes in binary systems, the value of the black hole spin parameter is normally determined by the formation mechanism, because the amount of mass that can be transferred from the companion star to the black hole is modest and cannot appreciably change the black hole mass and spin angular momentum [50]. In the case of supermassive black holes in galactic nuclei, the black hole spin parameter should be determined by the mass transferred from the host environment to the compact object and spin measurements can test different scenarios of galaxy evolution and merger [51, 52].
@@
A crucial ingredient in most black hole spin measurements with X-ray observations is the assumption that the inner edge of the accretion disk is at the innermost stable circular orbit (ISCO) of the Kerr spacetime [53]. There is indeed a one-to-one correspondence between the ISCO radius $R_{\rm ISCO}$ and the black hole spin parameter $a_{*}$ (see Fig. 3), so a precise and accurate measurement of the location of the inner edge of the accretion disk can be converted to a measurement of the black hole spin parameter. The assumption that the inner edge of the accretion disk is at the ISCO radius is well supported by observational and theoretical studies for black holes with an Eddington luminosity between a few percent and about 30% and a spectrum dominated by the thermal component of the accretion disk [54, 55]. It is still a controversial issue if this is also the case for sources in which the spectrum is dominated by the Comptonized spectrum from the corona [56]: the disk is probably truncated at a radius larger than the ISCO at low luminosities and the inner edge of the disk moves to the ISCO radius as the luminosity increases [57, 58].
@@
Figure 3: ISCO radius $R_{\rm ISCO}$ in the Kerr spacetime in Boyer-Lindquist coordinates as a function of the black hole spin parameter $a_{*}$. The lower (upper) curve is for co-rotating (counter-rotating) orbits, namely for orbits with angular momentum parallel (anti-parallel) to the black hole spin. $R_{\rm ISCO}$ in units $M=1$.
@@
Although the predictions of General Relativity have been extensively tested in weak gravitational fields with Solar System experiments and radio observations of binary pulsars [59], the strong-field regime is still largely unexplored. Black holes are ideal laboratories for testing General Relativity in the strong field regime as they are the sources of the strongest gravitational fields that can be found today in the Universe [60, 61]. The past 10 years have seen significant advancements in our capability of testing strong gravity and today we can test the nature of astrophysical black holes with X-ray, gravitational waves, and black hole imaging [62, 63, 64, 65]. Current constraints are still weak (especially if compared to our capability of testing atomic and particle physics) but they can be improved with the next generation of observational facilities.
2.1 Mass measurements
@@ 2.1.1 Spectroscopic mass estimates
In black hole X-ray binaries (XRBs) the physical estimate of radius in length units can be obtained using thermal continuum spectroscopy, where the accretion disk’s multi-temperature blackbody spectrum is fitted, together with an estimate for the inclination of the inner disk and the distance to the source. From the Stefan-Boltzmann law and appropriate modifications for the disk atmosphere, the emitting area and hence inner radius of the disk can be determined (see e.g. [66, 67] and Section 2.2.3). Alternatively, the measurement of physical radius can be carried out with X-ray polarimetric measurements of the disk thermal spectrum ([68, 69, 70] and Section 2.2.2).
@@
2.1.2 X-ray reverberation mapping estimates of black holes mass
@@
Figure 4: Comparison between eXTP-SFA (left panel) and XMM-Newton epic-PN (right panel) of the $\Delta\chi^{2}$ 2D contour plots between the black hole mass and the height of the lamppost corona. Both panels are the result of fitting simultaneously the time-averaged and lag energy spectra accounting for hard continuum lags and X-ray reverberation lags. The simulations are performed considering an AGN of $10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$ and an illuminating corona placed at 6 Rg above the black hole axis, the target is observed for 500 ks at 1 mCrab.
@@ 2.1.3 Mass estimates from X-ray variability and QPOs
Measurements of the X-ray variability of AGNs can provide constraints on the central black hole mass, since the amplitude of short-term ($<1$ day) X-ray flux variability anti-correlates with the black hole mass. As the metric of flux variability, the “normalized excess variance” $\sigma^{2}_{\text{rms}}$ can be calculated using the following equation,
@@
where $N$ is the number of observational epochs, $X_{i}$ is the individual X-ray flux measurement with $\sigma_{i,\textrm{err}}$ as its uncertainty, and $\mu$ is the mean flux value.
The empirical relation between $\sigma^{2}_{\text{rms}}$ and black hole mass has been calibrated. With $\sigma^{2}_{\text{rms}}$ calculated, the black hole mass can be constrained. With its long ($>100$ ks) uninterrupted exposures and excellent SFA collecting area and sensitivity, eXTP will extend the measurement of masses using the excess variance to much fainter AGN than were previously accessible. In particular, the method can be applied to many AGN with relatively low black hole masses and modest accretion rates, to check the scaling of the relation from higher-mass black holes observed in more luminous AGN.
@@ 2.2.1 X-ray reflection spectroscopy
Figure 5: The recoverability of the black hole spin parameter for eXTP/SFA and NICER/XTI with simulated black hole X-ray binary reflection spectra. The relxill model is used assuming a photon index $\Gamma=1.7$ and cutoff energy $E_{\rm cut}=300$ keV. The accretion disk extends to the ISCO and is viewed at an inclination angle of 45 degrees. Other parameters are the reflection fraction ($R_{f}$=1), the emissivity profile ($q=3$), the iron abundance ($A_{\rm Fe}=1$), the ionization parameter ($\log\xi=3$) and the flux ($2\times 10^{-8}$ erg cm-2 s-1 in the 2–10 keV band). We fit the simulated data in the range 4-10 keV. The dashed line represents the condition where the input spin for the simulation is equal to the best-fit spin ($a_{\rm in}=a_{\rm fit}$). The blue and red data points correspond to the best-fit values for NICER/XTI and eXTP/SFA, respectively. The shaded areas represent the 90% confidence intervals.
@@
The recoverability of the black hole spin parameter using X-ray reflection spectroscopy is shown in Figure 5 for eXTP/SFA, in comparison to NICER/XTI. We simulated 30 ks observations of bright black hole X-ray binaries (flux $2\times 10^{-8}$ erg cm-2 s-1 in the 2-10 keV energy band) with a black hole spin parameter $a_{*}$ ranging from 0 to 0.975. For the simulations, we used the publicly available model relxill [78, 79], assuming that the coronal spectrum has photon index $\Gamma=1.7$ and high-energy cutoff $E_{\rm cut}=300$ keV, the reflection fraction is $R_{\rm f}=1$, the inclination angle of the disk is $i=45$ deg, the inner edge of the accretion disk is at the ISCO radius of the spacetime, the emissivity profile of the disk is described by a simple power law with emissivity index $q=3$, the iron abundance of the disk is the Solar iron abundance ($A_{\rm Fe}=1$), and the ionization parameter of the disk is $\log\xi=3$ ($\xi$ in units erg cm s-1). We fit the simulated data in the range 4-10 keV, as current reflection models are unsuitable to fit the data of black hole X-ray binaries below 4 keV because of the high temperatures of their accretion disks. In general, the spin is better constrained if the black hole is rotating faster. From Fig. 5, it is also clear that eXTP/SFA can measure the black hole spin parameter with smaller uncertainties than NICER.
@@
Figure 6: The recoverability of corona inclinations for eXTP/SFA (red) and NICER/XTI (blue) with simulated black hole X-ray binary reflection spectra. The black hole spin is assumed to $a_{*}=0.998$. The radius and azimuth angle of the corona are assumed to $r_{c}=6\,R_{g}$ and $\phi_{c}=0^{\circ}$, respectively, and the corona moves radially outward at a speed $v_{r}=0.5\,c$. The values of other parameters are the same as those in Figure 5. The dashed line represents that the input inclination for the simulation is equal to its best-fit value. The blue and red data points correspond to the best-fit values for NICER/XTI and eXTP/SFA, respectively. The shaded areas represent the 90% confidence intervals.
@@
If the geometry of the X-ray corona is different from the lamp-post assumption, the non-axisymmetric illumination of the accretion disk from the off-axis corona has effects on the reflection spectra. We simulated the spectrum provided by a corona away from the black hole spin axis using offaxxill model [80, 81], which is an extension of relxill and allows the corona to be located off-axis and to move with arbitrary velocity. Figure 6 shows the recoverability of corona inclinations using X-ray reflection spectroscopy for eXTP/SFA. The black hole spin is assumed to $a_{*}=0.998$ in simulations. The radius and azimuth angle of the corona are assumed to $r_{c}=6\,R_{g}$ and $\phi_{c}=0^{\circ}$, respectively, and the corona moves radially outward at a relativistic speed $v_{r}=0.5\,c$. The corona inclination is better constrained if the corona has a low or high inclination. It is constrained worse if the corona has a moderate inclination. Figure 6 shows that eXTP/SFA provides better constraints on corona inclinations by contrast.

### v2

@@ 1 Introduction
The analysis of the X-ray radiation emitted from the strong gravity region around black holes is a powerful tool to study the accretion process of the innermost region of the accretion disk, measure black hole masses and spins, and test Einstein’s theory of General Relativity (GR) in the strong field regime. The prototype of astrophysical system for X-ray observations is shown in the left panel in Fig. 1 and is normally referred to as the disk-corona model [6, 7]. The central black hole can be either a stellar-mass black hole in an X-ray binary (BHXRB) system or a supermassive black hole in an active galactic nucleus (AGN). The crucial ingredient is that the accretion disk is Keplerian, geometrically thin or ‘slim’, and optically thick. Such an accretion disk is cold because it can efficiently emit radiation [^1] Thin accretion disks are cold with respect to other accretion disk models (e.g., advection-dominated accretion flow or ADAF [8]), as well as with respect to the corona. The thermal spectrum of the accretion disk turns out to be peaked in the soft X-ray band for stellar-mass black holes in X-ray binary systems and in the UV band for supermassive black holes in active galactic nuclei [9, 10]. The “corona” is some hot plasma ($\sim 100$ keV) near the black hole and the inner part of the accretion disk, but its exact morphology is not yet completely understood [11, 12, 13, 14, 15, 16, 17]. The corona may be the base of the jet, the hot atmosphere above the accretion disk, an inner, advection-dominated accretion flow when the inner edge of the disk is truncated at large radii, etc. (see the right panel in Fig. 1).
@@
    Are astrophysical black holes the Kerr black holes predicted by GR? Do astrophysical black holes have an event horizon or are they instead compact horizonless objects as predicted by a number of scenarios beyond GR [28]?
@@
    What is the mechanism responsible for QPOs? Are they generated in the accretion disk or in the corona? Can we use QPO observations for measuring black hole masses and spins and for testing Einstein’s theory of GR?
@@
The enhanced X-ray Timing and Polarimetry (eXTP) mission has been proposed by a consortium led by the Institute of High-Energy Physics of the Chinese Academy of Sciences to address these questions. Its launch is currently scheduled in 2030. The primary goals of eXTP are: $i)$ the study of matter at very high density [29], ii)ii) the study of strong gravity with black holes (this paper), $iii)$ the study of matter in strong magnetic fields [30], $iv)$ the study of violent high-energy astrophysical phenomena with X-ray in the era of multi-messanger astrophysics [31] and $v$ the observatory science [32].
@@
The W2C is a secondary instrument of the science payload, featuring a coded mask camera with a FoV of approximately 1500 square degrees (Full-Width Zero Response, FWZR). The instrument achieves a sensitivity of $4\times 10^{-7}\,{\rm erg\,cm^{-2}\,s^{-1}}$ (1 s exposure) across the 10–600 keV energy range, with an angular resolution of $20^{\prime}$, a position accuracy of 5’, and an energy resolution better than $30\%$ at 60 keV.
@@
Beyond all-sky monitoring, eXTP will conduct detailed studies of X-ray novae in the future, with a particular focus on black hole X-ray binaries in this paper. The top panels of Fig. 2 showcase a simulated SFA light curve and spectrum of the black hole X-ray binary GRS 1915+105 [40], based on archival observations from Insight-HXMT. eXTP will reveal rapid and periodic variability on short timescales associated with disk instability [see Section 2.2 or reference 41]. A 20 ks eXTP observation will also demonstrate the spectral capabilities of SFA, detecting relativistic Fe K disk emission as part of disk reflection (discussed in Section 2.2.1), which will provide insight into the properties of the strong gravity region and the accretion disk. Additionally, eXTP will resolve the absorption features of the disk winds, allowing measurements of their velocities and helping to determine their kinematic properties and potential launching mechanisms (see Section 3.1.2).
Figure 2: The top two panels show a simulated SFA light curve and spectrum of GRS 1915++105, demonstrating eXTP’s capabilities in studying strong-gravity accretion physics. The rapid, periodic variability reveals instabilities of the inner accretion disk of this black hole; the relativistic Fe K emission line provides insights into strong-gravity effects and accretion disk properties; and the absorption features from disk winds enable kinematic studies and an investigation of potential driving mechanisms. These topics will be further explored throughout this paper. The bottom two panels illustrate the expected number of black hole X-ray novae eXTP will observe over the first three-year period, based on historical events. These include short outbursts lasting a few months and longer outbursts extending over several years. See text for further details.
The paper is organized as follows. In Section 2, we show how eXTP can improve current measurements of black hole masses and spins as well as put stronger constraints on possible deviations from the predictions of GR. In Section 3, we show how eXTP can help us to understand better the accretion process in the strong gravity region of black holes. Summary and conclusions are reported in Section 4.
2 Measuring the Basic Parameters of Black Holes and Testing GR
@@
For a Kerr black hole, the mass $M$ is a free parameter, and therefore we may have black holes with arbitrarily low as well as arbitrarily high masses. However, the black holes in our Universe must form through some astrophysical process and therefore it is natural to expect black holes with masses only in specific ranges. This is exactly what we find. Astrophysical observations have discovered at least two classes of black holes: stellar-mass black holes ($\approx 3-100 MM_{\odot}$[46] and supermassive black holes ($\sim 10^{5}-10^{10} MM_{\odot}$[47]. There is likely a third class of objects, the so-called intermediate mass black hole candidates [48], which are black hole candidates with a mass filling the gap between the stellar-mass and the supermassive ones.
@@
Unlike the mass, the spin angular momentum $J$ of a Kerr black hole is subject to certain constraints. If we define the dimensionless spin parameter $a_{*}=J/M^{2}$ (in units in which $c=G_{\rm N}=1$), the condition for the existence of a Kerr black hole is $|a_{*}|\leq 1$ (Kerr bound). For $|a_{*}|>1$, there is no event horizon and the Kerr solution describes the spacetime of a naked singularity, which is thought to be impossible to create in the Universe [49]. Spin measurements are crucial to understand how black holes form, evolve, and interact with the host environment [50]. In the case of stellar-mass black holes in binary systems, it is often thought that the value of the black hole spin parameter is determined by the formation mechanism (and the natal spin is expected to be low [51, 52, 53]), because the amount of mass that can be transferred from the companion star to the black hole is modest and cannot appreciably change the black hole mass and spin angular momentum [54, 55, 56]. However, such a conclusion is not necessarily true [57]. In the case of low-mass XRBs, the black hole may be spun up immediately after its formation [58]. In the case of high-mass XRBs, the black hole spin may be directly related to the angular momentum of the progenitor star, which can transfer part of its envelope to the companion star and get a core with high angular momentum [59]. See Ref. [60] for a recent review on this topic. In the case of supermassive black holes in galactic nuclei, the black hole spin parameter should be determined by the mass transferred from the host environment to the compact object and spin measurements can test different scenarios of galaxy evolution and merger [61, 62].
@@
A crucial ingredient in most black hole spin measurements with X-ray observations is the assumption that the inner edge of the accretion disk is at the innermost stable circular orbit (ISCO) of the Kerr spacetime [63]. There is indeed a one-to-one correspondence between the ISCO radius $R_{\rm ISCO}$ and the black hole spin parameter $a_{*}$ (see Fig. 3), so a precise and accurate measurement of the location of the inner edge of the accretion disk can be converted to a measurement of the black hole spin parameter. The assumption that the inner edge of the accretion disk is at the ISCO radius is well supported by observational and theoretical studies for black holes with an Eddington luminosity between a few percent and about 30% and a spectrum dominated by the thermal component of the accretion disk (soft state) [64, 65]. It is still a controversial issue if this is also the case for sources in which the spectrum is dominated by the Comptonized spectrum from the corona (hard state) [66] [^2] We note that both hard and soft states are observed at luminosities $1\%$ of the Eddington luminosity $L_{\rm Edd}$, implying that both corresponding physical solutions can exist at similar accretion rates.: the disk is probably truncated at a radius larger than the ISCO at low luminosities and the inner edge of the disk moves to the ISCO radius as the luminosity increases [67, 68] (see, however, Refs. [69, 70]).
@@
Figure 3: ISCO radius $R_{\rm ISCO}$ in the Kerr spacetime in Boyer-Lindquist coordinates as a function of the black hole spin parameter $a_{*}$. $a_{*}>0$ ($a_{*}<0$) is for co-rotating (counter-rotating) orbits, namely for orbits with angular momentum parallel (anti-parallel) to the black hole spin. $R_{\rm ISCO}$ in units $M=1$.
@@
Although the predictions of GR have been extensively tested in weak gravitational fields with Solar System experiments and radio observations of binary pulsars [71], the strong-field regime is still largely unexplored. Black holes are ideal laboratories for testing GR in the strong field regime as they are the sources of the strongest gravitational fields that can be found today in the Universe [72, 73]. The past 10 years have seen significant advancements in our capability of testing strong gravity and today we can test the nature of astrophysical black holes with X-ray, gravitational waves, and black hole imaging [74, 75, 76, 77]. Current constraints are still weak (especially if compared to our capability of testing atomic and particle physics) but they can be improved with the next generation of observational facilities.
2.1 Mass measurements in X-ray binaries
@@ 2.1.1 Spectroscopic mass estimates
In black hole X-ray binaries (BHXRBs) the physical estimate of radius in length units can be obtained using thermal continuum spectroscopy, where the accretion disk’s multi-temperature blackbody spectrum is fitted, together with an estimate for the inclination of the inner disk and the distance to the source. From the Stefan-Boltzmann law and appropriate modifications for the disk atmosphere, the emitting area and hence inner radius of the disk can be determined (see e.g. [78, 79] and Section 2.2.3). Alternatively, the measurement of physical radius can be carried out with X-ray polarimetric measurements of the disk thermal spectrum ([80, 81, 82] and Section 2.2.2).
@@
2.1.2 X-ray reverberation mapping estimates of black hole mass
@@
Figure 4: Comparison between eXTP-SFA (left panel) and XMM-Newton epic-PN (right panel) of the $\Delta\chi^{2}$ 2D contour plots between the black hole mass and the height of the lamp-post corona. Both panels are the result of fitting simultaneously the time-averaged and lag energy spectra accounting for hard continuum lags and X-ray reverberation lags. The simulations are performed considering an AGN of $10^{7}\leavevmode\nobreak\ M_{\odot}$ and an illuminating corona placed at 6 $R_{g}$ above the black hole axis, the target is observed for 500 ks at 1 mCrab.
@@ 2.1.3 Mass estimates from X-ray variability and QPOs
Measurements of the X-ray variability of AGNs can provide constraints on the central black hole mass, since the amplitude of short-term (<1<1 day) X-ray flux variability anti-correlates with the black hole mass. As the metric of flux variability, the “normalized excess variance” $\sigma^{2}_{\text{rms}}$ is defined as the following equation,
@@
where NN is the number of observational epochs, $X_{i}$ is the individual X-ray flux measurement with $\sigma_{i,\textrm{err}}$ as its uncertainty, and $\mu$ is the mean flux value.
The empirical relation between $\sigma^{2}_{\text{rms}}$ and black hole mass has been calibrated (e.g. [88]). With $\sigma^{2}_{\text{rms}}$ calculated, the black hole mass can be constrained. With its long (>100>100 ks) uninterrupted exposures and excellent SFA collecting area and sensitivity, eXTP will extend the measurement of masses using the excess variance to much fainter AGN than were previously accessible. In particular, the method can be applied to many AGN with relatively low black hole masses and modest accretion rates, to check the scaling of the relation from higher-mass black holes observed in more luminous AGN.
@@ 2.2.1 X-ray reflection spectroscopy
Figure 5: The recoverability of the black hole spin parameter for eXTP/SFA and NICER/XTI with simulated BHXRB reflection spectra. The relxill model is used assuming a photon index $\Gamma=1.7$ and cutoff energy $E_{\rm cut}=300$ keV. The accretion disk extends to the ISCO and is viewed at an inclination angle of 45 degrees. Other parameters are the reflection fraction ($R_{f}$=1), the emissivity profile ($q=3$), the iron abundance ($A_{\rm Fe}=1$), the ionization parameter ($\log\xi=3$) and the flux ($2\times 10^{-8}$ erg cm-2 s-1 in the 2–10 keV band). We fit the simulated data in the range 4-10 keV. The dashed line represents the condition where the input spin for the simulation is equal to the best-fit spin ($a_{\rm in}=a_{\rm fit}$). The blue and red data points correspond to the best-fit values for NICER/XTI and eXTP/SFA, respectively. The shaded areas represent the 90% confidence intervals.
@@
The recoverability of the black hole spin parameter using X-ray reflection spectroscopy is shown in Fig. 5 for eXTP/SFA, in comparison to NICER/XTI. We simulated 30 ks observations of bright black hole X-ray binaries (flux $2\times 10^{-8}$ erg cm-2 s-1 in the 2-10 keV energy band) with a black hole spin parameter $a_{*}$ ranging from 0 to 0.975. For the simulations, we used the publicly available model relxill [91, 92], assuming that the coronal spectrum has photon index $\Gamma=1.7$ and high-energy cutoff $E_{\rm cut}=300$ keV, the reflection fraction is $R_{\rm f}=1$, the inclination angle of the disk is $i=45$ deg, the inner edge of the accretion disk is at the ISCO radius of the spacetime, the emissivity profile of the disk is described by a simple power law with emissivity index $q=3$, the iron abundance of the disk is the Solar iron abundance ($A_{\rm Fe}=1$), and the ionization parameter of the disk is $\log\xi=3$ ($\xi$ in units erg cm s-1). We fit the simulated data in the range 4-10 keV, as current reflection models are unsuitable to fit the data of black hole X-ray binaries below 4 keV because of the high temperatures of their accretion disks. In general, the spin is better constrained if the black hole is rotating faster. From Fig. 5, it is also clear that eXTP/SFA can measure the black hole spin parameter with smaller uncertainties than NICER.
@@
Figure 6: The recoverability of corona inclinations for eXTP/SFA (red) and NICER/XTI (blue) with simulated black hole X-ray binary reflection spectra. The black hole spin is assumed to $a_{*}=0.998$. The radius and azimuth angle of the corona are assumed to $r_{c}=6\,R_{g}$ and $\phi_{c}=0^{\circ}$, respectively, and the corona moves radially outward at a speed $v_{r}=0.5\,c$. The values of other parameters are the same as those in Fig. 5. The dashed line represents that the input inclination for the simulation is equal to its best-fit value. The blue and red data points correspond to the best-fit values for NICER/XTI and eXTP/SFA, respectively. The shaded areas represent the 90% confidence intervals.
@@
For simplicity we initially assumed a lamp-post coronal geometry which is axisymmetric. Deviations from axisymmetry might be expected for precessing coronae (e.g. due to differential disk illumination in the precession cycle or by causing a misalignment of a coronal ’jet base’ with the thin disk), or in case parts of the disk are obscured relative to the observer (e.g. by the corona itself). Such situations are too complex to model directly for our purposes here, so we show the SFA sensitivity to such effects by considering non-axisymmetric illumination of the accretion disk from an off-axis lamp-post corona. We simulated the spectrum provided by a corona away from the black hole spin axis using offaxxill model [93, 94], which is an extension of relxill and allows the corona to be located off-axis and to move with arbitrary velocity. Fig. 6 shows the recoverability of corona inclinations using X-ray reflection spectroscopy for eXTP/SFA. The black hole spin is assumed to $a_{*}=0.998$ in simulations. The radius and azimuth angle of the corona are assumed to $r_{c}=6\,R_{g}$ and $\phi_{c}=0^{\circ}$, respectively, and the corona moves radially outward at a relativistic speed $v_{r}=0.5\,c$. The corona inclination is better constrained if the corona has a low or high inclination. It is constrained worse if the corona has a moderate inclination. Fig. 6 shows that eXTP/SFA provides better constraints on corona inclinations by contrast with NICER.